OpenStudy (anonymous):

If a, b and c are positive, and if ab= 2, bc = 3 and ac = 6, what is a + b + c?

OpenStudy (anonymous):

trial and error i have my answer which is 6, but i dont know the solution

OpenStudy (anonymous):

just substitute...

OpenStudy (anonymous):

b=2/a, c=6/a so from bc = 3, 2/a * 6/a = 3. so a=2. b=1, c=3... a+b+c = 6

OpenStudy (anonymous):

ab*bc*ac=6*6 abc = 6 a+b+c =x a(abc) +abc*b + abc*c = 6x 6*2 + 3*2 + 6*3 = 6x x = 2 + 1 + 3 = 6

OpenStudy (anonymous):

ohhh!!!

OpenStudy (anonymous):

@Ishaan94 , that's nice!!!

OpenStudy (anonymous):

Thanks :-)

OpenStudy (anonymous):

nice solutions .. thanks!

OpenStudy (anonymous):

hey @Ishaan94 and @dpalnc ab*bc*ca=6*6 abc=6 using AM>=GM (a+b+c)/3>=(abc)^(1/3) a+b+c>=3(6)^(1/3)=5.45... is it correct...??